For the reaction KOH (aq) + HCN (aq) → H2O (l) + KCN (aq), what volume of HCN at 2.00 M would be required in order to make 33.0 g of KCN, when there is more HCN than needed?

A. 25.4 L
B. 2.54 L
C. 0.254 L
D. 0.0254 L
E. None of the Above
1
Expert's answer
2018-04-05T07:36:10-0400
Answer
n(HCN) = m(KCN)*1 mol(HCN)/(1 mol(KCN)*M(KCN), if M(KCN) = 65 g/mol:
n(HCN) = 33*1/65*1 = 0,508 mol
Molarity(HCN) or c(HCN) = n(HCN)/V(HCN), if c(HCN) = 2M than:
V(HCN) = n(HCN)/c(HCN) = 0,508/2 = 0,254 L
Answer “C” is true.